# Surveying the MLLM Landscape: A Meta-Review of Current Surveys

### v1

@@ 5.2 Datasets and Training
The field is characterized by massive, multimodal datasets such as MS-COCO, Visual Genome, and custom-curated sets like LAION. Pretraining on large-scale datasets remains a core strategy, with some surveys, like “Large-scale Multi-Modal Pre-trained Models”, etc.[81], offering a comprehensive taxonomy of pretraining methodologies.
@@ 5.3 Evaluation and Metrics
Evaluation challenges are a key theme, where traditional language or vision metrics fall short. Cross-modal retrieval, image captioning, and visual question answering (VQA) are popular benchmarks, but new methods to evaluate multimodal coherence and reasoning are discussed in evaluation-focused surveys like “A Survey on Evaluation of Large Language Models”, etc.[57, 82, 83].
@@ 5.3.1 Security: Adversarial Attacks
Adversarial attacks have been a critical concern for MLLMs. Caffagni et al.[84] highlight that MLLMs are particularly susceptible to adversarial attacks that exploit weaknesses in visual inputs. By introducing imperceptible noise into images, attackers can manipulate the vision encoder, leading to incorrect or even harmful text generation. Wang et al.[85] elaborate on jailbreaking attacks, where adversaries bypass the model’s safety alignment mechanisms. Techniques like prompt injection can disrupt the model’s chain-of-thought reasoning, leading to the generation of dangerous or inappropriate content. Especially in white-box attacks, attackers utilize the model’s gradient information to craft adversarial examples that precisely control the model’s outputs.[86]
Zhao et al.[87] emphasize the complexity of multimodal adversarial attacks, where both image and text inputs are manipulated. These attacks exploit the model’s difficulty in handling cross-modal perturbations, making it challenging to detect malicious inputs. To counter these attacks, researchers suggest improving cross-modal alignment algorithms and developing new defense mechanisms to enhance robustness across modalities.
Another vulnerability arises in the cross-modal alignment process, which connects the features extracted from images with those from textual data. As Shayegani et al.[88] explain, attackers can exploit this process to disrupt the model’s understanding of multimodal data, compromising its predictions. This highlights a need for robust alignment mechanisms that can defend against such manipulations.
@@ 5.3.2 Bias: Hallucinations and Data Bias
Liu et al.[89] describe the hallucination phenomenon in Multimodal Large Language Models (MLLMs), where the generated outputs deviate from the visual input, leading to the fabrication of objects or misrepresentation of relationships in an image. This issue is often a byproduct of data bias in the training sets, as well as the limitations of vision encoders in accurately grounding images.
Hallucinations can be traced to several causes, ranging from the models’ reliance on noisy or incomplete training data to the inherent limitations of cross-modal alignment mechanisms[61]. For instance, models may incorporate statistically frequent objects that are not present in the specific image or confuse the relationships between depicted elements. This happens because MLLMs may memorize biases present in training datasets, such as common object pairings or overly simplistic image-text associations. As a result, hallucinations not only mislead users but also exacerbate existing biases in the models’ outputs, leading to incorrect assumptions about objects’ existence, attributes, or relationships[60].
Addressing these hallucinations requires improving the modality alignment mechanisms to better synchronize textual and visual data. This involves enhancing the capabilities of vision encoders to more accurately represent fine-grained details of the images and ensuring that the language models respect the constraints imposed by visual context. Bai et al.[61] advocate for the use of specialized evaluation benchmarks and techniques, such as leveraging better-aligned datasets or employing post-processing corrections. Furthermore, strategies like counterfactual data augmentation and the introduction of diverse visual instructions can help models generalize better to unseen scenarios, thus reducing the frequency of hallucinations and aligning model predictions with true visual inputs[90].
@@ 5.3.3 Fairness: Adversarial Training and Human Feedback
To address both security and fairness, several defense strategies have been proposed. Shayegani et al.[88] suggest employing adversarial training and data augmentation as methods to strengthen the model’s robustness. By introducing adversarial examples during the training phase, models can learn to identify and resist adversarial inputs. Furthermore, safety steering vectors and multimodal input validation can dynamically detect and correct potentially biased or adversarial outputs during inference.
Liu et al.[91] explore techniques to improve fairness by leveraging Reinforcement Learning from Human Feedback (RLHF). This method allows models to adjust their outputs based on user feedback, ensuring that the generated text aligns with societal values and ethical standards. The approach ensures that models maintain fairness by incorporating diverse viewpoints and mitigating biases present in training data.
@@ 5.3.4 Defense Against Jailbreaking Attacks
Jin et al.[92] classify different types of jailbreaking attacks and propose defense strategies such as prompt tuning and gradient-based defenses. These approaches focus on adjusting the input prompts and strengthening alignment algorithms to reduce the success rate of jailbreaking attacks, ensuring that the model remains secure and aligned with ethical guidelines.
@@ 6.1 Current Trends
One of the most prominent trends in the surveyed literature is the enhanced integration of multiple modalities—such as text, images, and audio—within MLLMs. Surveys like “The (R)Evolution of Multimodal Large Language Models”, etc.[93, 94, 95, 96] emphasize the shift from unimodal to multimodal systems as a transformative leap, enabling models to more closely mimic human perception. This trend is echoed across multiple papers, highlighting the field’s focus on achieving a more holistic understanding of information by combining different types of data inputs.
@@
A recurring theme in recent surveys is the expansion of MLLM applications into various domains, particularly those requiring complex, multimodal understanding, such as autonomous agents and medical diagnostics. For instance, the survey “A Survey on Multimodal Large Language Models”[97] traces the significant impact of MLLMs across different industries, suggesting a growing trend toward domain-specific adaptations of these models. This trend points to the increasing specialization of MLLMs to meet the demands of specific fields.
@@
As MLLMs become more prevalent, the need for robust evaluation metrics and benchmarking has become increasingly critical. The surveyed literature frequently discusses the development of new benchmarks designed to assess the performance of MLLMs across different modalities. For example, “A Survey of Large Language Models”[98] provides an in-depth analysis of existing evaluation methodologies, indicating a trend towards more comprehensive and standardized performance assessments.
@@
These trends underscore the rapid evolution of MLLMs and the diverse challenges and opportunities they present. The surveyed literature not only identifies these key areas but also sets the stage for future research directions, which are discussed in the subsequent sections. [99, 100]
@@ 6.2 Research Gaps
While most surveys focus on the integration of text, images, and, to some extent, audio, there is a noticeable gap in exploring the potential of other modalities, such as haptic feedback, olfactory data, and advanced sensory inputs. The survey like "The (R)Evolution of Multimodal Large Language Models", etc.[93, 101, 102, 103, 104] acknowledges the importance of holistic integration but primarily emphasizes traditional modalities. Future research could expand on how these lesser-known modalities can be incorporated into MLLMs to create even more comprehensive models.
@@
Current surveys tend to focus on the performance of MLLMs at a specific point in time, often lacking longitudinal studies that track the evolution of model capabilities and limitations over extended periods. For instance, "A Survey of Large Language Models"[98, 105, 106, 107] provides a snapshot of the state of MLLMs but does not address how these models might evolve with new data, architectures, or training techniques. Longitudinal studies could provide valuable insights into the long-term viability and scalability of MLLMs.
@@
There is limited coverage on the application of MLLMs across vastly different domains and the effectiveness of transfer learning in these contexts. Surveys such as "Large-scale Multi-Modal Pre-trained Models: A Survey"[81] touch upon domain-specific applications but do not fully explore how models trained in one domain perform when transferred to another. This is a critical area for research, especially in understanding the generalization capabilities of MLLMs across diverse fields.
@@
The intersection of multilingualism and multimodality is another underexplored area. Although "Multilingual Large Language Model: A Survey of Surveys"[108] addresses multilingual capabilities, it does not fully explore how these capabilities interact with multimodal inputs. Research in this area could lead to more inclusive MLLMs that better serve global populations by effectively integrating multilingual and multimodal data.

### v2

@@ 5.2 Datasets and Training
The field is characterized by massive, multimodal datasets such as MS-COCO, Visual Genome, and custom-curated sets like LAION. Pretraining on large-scale datasets remains a core strategy, with some surveys, like “Large-scale Multi-Modal Pre-trained Models”, etc.[wang2024largescalemultimodalpretrainedmodels], offering a comprehensive taxonomy of pretraining methodologies.
@@ 5.3 Evaluation and Metrics
Evaluation challenges are a key theme, where traditional language or vision metrics fall short. Cross-modal retrieval, image captioning, and visual question answering (VQA) are popular benchmarks, but new methods to evaluate multimodal coherence and reasoning are discussed in evaluation-focused surveys like “A Survey on Evaluation of Large Language Models”, etc.[chang2024survey, ge2024mllmbenchevaluatingmultimodalllms, chen2024mllmasajudgeassessingmultimodalllmasajudge].
@@ 5.3.1 Security: Adversarial Attacks
Adversarial attacks have been a critical concern for MLLMs. Caffagni et al.[caffagni2024revolutionmultimodallargelanguage] highlight that MLLMs are particularly susceptible to adversarial attacks that exploit weaknesses in visual inputs. By introducing imperceptible noise into images, attackers can manipulate the vision encoder, leading to incorrect or even harmful text generation. Wang et al.[wang2024llms] elaborate on jailbreaking attacks, where adversaries bypass the model’s safety alignment mechanisms. Techniques like prompt injection can disrupt the model’s chain-of-thought reasoning, leading to the generation of dangerous or inappropriate content. Especially in white-box attacks, attackers utilize the model’s gradient information to craft adversarial examples that precisely control the model’s outputs.[pi2024mllmprotectorensuringmllmssafety]
Zhao et al.[zhao2024survey] emphasize the complexity of multimodal adversarial attacks, where both image and text inputs are manipulated. These attacks exploit the model’s difficulty in handling cross-modal perturbations, making it challenging to detect malicious inputs. To counter these attacks, researchers suggest improving cross-modal alignment algorithms and developing new defense mechanisms to enhance robustness across modalities.
Another vulnerability arises in the cross-modal alignment process, which connects the features extracted from images with those from textual data. As Shayegani et al.[shayegani2023survey] explain, attackers can exploit this process to disrupt the model’s understanding of multimodal data, compromising its predictions. This highlights a need for robust alignment mechanisms that can defend against such manipulations.
@@ 5.3.2 Bias: Hallucinations and Data Bias
Liu et al.[liu2024surveyhallucinationlargevisionlanguage] describe the hallucination phenomenon in Multimodal Large Language Models (MLLMs), where the generated outputs deviate from the visual input, leading to the fabrication of objects or misrepresentation of relationships in an image. This issue is often a byproduct of data bias in the training sets, as well as the limitations of vision encoders in accurately grounding images.
Hallucinations can be traced to several causes, ranging from the models’ reliance on noisy or incomplete training data to the inherent limitations of cross-modal alignment mechanisms[bai2024hallucination]. For instance, models may incorporate statistically frequent objects that are not present in the specific image or confuse the relationships between depicted elements. This happens because MLLMs may memorize biases present in training datasets, such as common object pairings or overly simplistic image-text associations. As a result, hallucinations not only mislead users but also exacerbate existing biases in the models’ outputs, leading to incorrect assumptions about objects’ existence, attributes, or relationships[gallegos2024bias].
Addressing these hallucinations requires improving the modality alignment mechanisms to better synchronize textual and visual data. This involves enhancing the capabilities of vision encoders to more accurately represent fine-grained details of the images and ensuring that the language models respect the constraints imposed by visual context. Bai et al.[bai2024hallucination] advocate for the use of specialized evaluation benchmarks and techniques, such as leveraging better-aligned datasets or employing post-processing corrections. Furthermore, strategies like counterfactual data augmentation and the introduction of diverse visual instructions can help models generalize better to unseen scenarios, thus reducing the frequency of hallucinations and aligning model predictions with true visual inputs[liang2024internal].
@@ 5.3.3 Fairness: Adversarial Training and Human Feedback
To address both security and fairness, several defense strategies have been proposed. Shayegani et al.[shayegani2023survey] suggest employing adversarial training and data augmentation as methods to strengthen the model’s robustness. By introducing adversarial examples during the training phase, models can learn to identify and resist adversarial inputs. Furthermore, safety steering vectors and multimodal input validation can dynamically detect and correct potentially biased or adversarial outputs during inference.
Liu et al.[liu2023trustworthy] explore techniques to improve fairness by leveraging Reinforcement Learning from Human Feedback (RLHF). This method allows models to adjust their outputs based on user feedback, ensuring that the generated text aligns with societal values and ethical standards. The approach ensures that models maintain fairness by incorporating diverse viewpoints and mitigating biases present in training data.
@@ 5.3.4 Defense Against Jailbreaking Attacks
Jin et al.[jin2024jailbreakzoo] classify different types of jailbreaking attacks and propose defense strategies such as prompt tuning and gradient-based defenses. These approaches focus on adjusting the input prompts and strengthening alignment algorithms to reduce the success rate of jailbreaking attacks, ensuring that the model remains secure and aligned with ethical guidelines.
@@ 6.1 Current Trends
One of the most prominent trends in the surveyed literature is the enhanced integration of multiple modalities—such as text, images, and audio—within MLLMs. Surveys like “The (R)Evolution of Multimodal Large Language Models”, etc.[caffagni2024r, hamadi2023large, song2023bridge, zhou2023comprehensive] emphasize the shift from unimodal to multimodal systems as a transformative leap, enabling models to more closely mimic human perception. This trend is echoed across multiple papers, highlighting the field’s focus on achieving a more holistic understanding of information by combining different types of data inputs.
@@
A recurring theme in recent surveys is the expansion of MLLM applications into various domains, particularly those requiring complex, multimodal understanding, such as autonomous agents and medical diagnostics. For instance, the survey “A Survey on Multimodal Large Language Models”[yin2023survey] traces the significant impact of MLLMs across different industries, suggesting a growing trend toward domain-specific adaptations of these models. This trend points to the increasing specialization of MLLMs to meet the demands of specific fields.
@@
As MLLMs become more prevalent, the need for robust evaluation metrics and benchmarking has become increasingly critical. The surveyed literature frequently discusses the development of new benchmarks designed to assess the performance of MLLMs across different modalities. For example, “A Survey of Large Language Models”[zhao2023surveylargelanguagemodels] provides an in-depth analysis of existing evaluation methodologies, indicating a trend towards more comprehensive and standardized performance assessments.
@@
These trends underscore the rapid evolution of MLLMs and the diverse challenges and opportunities they present. The surveyed literature not only identifies these key areas but also sets the stage for future research directions, which are discussed in the subsequent sections. [yao2024minicpmvgpt4vlevelmllm, ma2024eemllmdataefficientcomputeefficientmultimodal]
@@ 6.2 Research Gaps
While most surveys focus on the integration of text, images, and, to some extent, audio, there is a noticeable gap in exploring the potential of other modalities, such as haptic feedback, olfactory data, and advanced sensory inputs. The survey like "The (R)Evolution of Multimodal Large Language Models", etc.[caffagni2024r, hadi2023large, 10386743, awais2023foundational, yang2024harnessing] acknowledges the importance of holistic integration but primarily emphasizes traditional modalities. Future research could expand on how these lesser-known modalities can be incorporated into MLLMs to create even more comprehensive models.
@@
Current surveys tend to focus on the performance of MLLMs at a specific point in time, often lacking longitudinal studies that track the evolution of model capabilities and limitations over extended periods. For instance, "A Survey of Large Language Models"[zhao2023surveylargelanguagemodels, sohail2023future, kalyan2023survey, carolan2024review] provides a snapshot of the state of MLLMs but does not address how these models might evolve with new data, architectures, or training techniques. Longitudinal studies could provide valuable insights into the long-term viability and scalability of MLLMs.
@@
There is limited coverage on the application of MLLMs across vastly different domains and the effectiveness of transfer learning in these contexts. Surveys such as "Large-scale Multi-Modal Pre-trained Models: A Survey"[wang2024largescalemultimodalpretrainedmodels] touch upon domain-specific applications but do not fully explore how models trained in one domain perform when transferred to another. This is a critical area for research, especially in understanding the generalization capabilities of MLLMs across diverse fields.
@@
The intersection of multilingualism and multimodality is another underexplored area. Although "Multilingual Large Language Model: A Survey of Surveys"[qin2024multilinguallargelanguagemodel] addresses multilingual capabilities, it does not fully explore how these capabilities interact with multimodal inputs. Research in this area could lead to more inclusive MLLMs that better serve global populations by effectively integrating multilingual and multimodal data.
@@ 7.2 Future Directions
Figure 2: Distribution of Survey Domains in MLLM Studies
Figure 3: Word Cloud Of Survey Titles In MLLM Studies